# SF-MMCN:A Low Power Re-configurable Server Flow Convolution Neural Network Accelerator

### v1

@@ 3.5 Data Reuse
Leveraging the SF structure, proposed SF-MMCN optimizes data reuse due to the prevalence of repeated input images, Figure 12 (a). Between each convolution cycle, there exist 8 repeated input data, corresponding to 8 registers from PE_1 to PE_8. To accommodate the 16-bit requirements for residual blocks in hardware, SF-MMCN expands the bit-width of these 8 registers to 32 bits, facilitating simultaneous storage of residual and reused data, Figure 12 (b). This strategic approach reduces power consumption by avoiding the reloading of repeated data in every computation cycle. As a result, SF-MMCN significantly conserves energy, demonstrating the efficacy of its SF structure.
@@
Figure 12: (a) Illustration of repeated data in proposed SF-MMCN. (b) Data-reused structure.
3.6 Implementation Architecture
@@
The implementation of proposed SF-MMCN is indicated in Figure 13. The input image features and weight data are stored in off-chip memory before executing. After input and weight data enter the computation core through I/O interface, dataflow would be managed by TOP CTRL through weight and input buffer.
Figure 14: Dataflow comparison between traditional structures and proposed SF-MMCN on (a) waveform. (b) CNN structure illustration.
There are four SF-MMCNs in implementation of proposed SF-MMCN. Owing to data reuse in implementation and large MAC in dense layer, each SF-MMCN can exchange data by registers of each PE in SF-MMCN. If the output channel exceeds hardware limitation, the dataflow of computation will iterate to complete convolutions. What’s more, with a pooling unit and activation function unit in proposed SF-MMCN, proposed also support different functions of CNN as MMCN in Hsu et al. (2023). All control signals are switched by TOP CTRL through testbench.
3.7 Operation Efficiency of Residual Blocks
In traditional CNN accelerators, when facing a parallel CNN structure, most of strategy is modifying the dataflow from parallel to series Mo et al. (2021), Hsu et al. (2023), as shows in Figure 14 (a). In this study, proposed SF-MMCN combine normal convolution layer and a convolution layer hide in residual block. The Residual_0 in figure 14 (a) doesn’t have to wait for series structure (Conv_0 and Conv_1 in Figure 14 (a)) finishing computing. In other words, according to Figure 14 (b), traditional CNN accelerators spend additional cycles to complete convolution layers with parallel structures. Therefore, proposed SF-MMCN perform the higher efficiency while conducting parallel structures.
3.8 Hardware Utilization and Power Efficiency
This paper aim to increases the utilization of PEs in PE array in a CNN accelerator. Due to pipeline technique in a PE in proposed SF-MMCN, single PE can complete a convolution computation by itself. The utilization of PE in hardware is indicated in:
@@ 3.8 Hardware Utilization and Power Efficiency
In equation (1), Where $C_{t}$ is the percentage of computing cycles, which means actual operating cycles $T$ divided by total enable cycles $t$. $PE_{act}$ and $PE_{total}$ represent actual executing PEs and total PEs in hardware respectively. $U_{PE}$ is the utilization of PEs in hardware. Although most CNN accelerators perform high PE utilization, the occurrence of redundant PE is still happened when encountering un-suitable CNN model. The utilization of PE is relatively connected with power consumption. As the biggest structure in a CNN accelerator, PE array consumes the most power in operation. The equation (3) indicates the total power $P_{total}$ of a CNN accelerator.
@@
N is equal to $PE_{act}$ in equation (2), which means the number of actual executing PEs in hardware. $P_{1}$ represents the power consumption of single PE, $P_{R}$ and $P_{C}$ are the power of redundant circuit and control unit respectively. However, given to the diversity of conditions in different CNN models, such as parallel block mentioned before. What’s more, with the development of CC accelerators, there are much more studies perform high efficiency in implementation. $P_{total}$ only expresses the estimation of total power. Therefore, another concept of efficiency factor ($\nu$) is explained in equation (4).
@@
Efficiency factor equals to the utilization of PE divided by total power consumption. According to equation (4), it reveals that if the value of $\nu$ increases, which means corresponded structure spend much more power on redundant hardware. This is because the value of PE utilization never equals to zero except the circuit is idle. Consequently, the smaller $\nu$ shows that the power consumption mostly result from PE array. Likewise, higher PE utilization represents the well allocated of hardware, which indicates more PEs are conducting MAC operations. In other words, to reduce the power consumption of a CNN accelerator, it’s necessary to reduce the scale of PE array, which meets the target of this paper.
Moreover, function (4) can also compute under single convolution layers. This paper emphasizes the operation conditions of series and parallel CNN models, function (4) also provides $\nu$ under different environments by adjusting the value of $N$ ans $P_{R}$ in equation (4).
4 Experimental results
Proposed SF-MMCN is realized in Verilog HDL and logic synthesis with Design Compiler by Synopsys and implemented in TSMC 40nm technology under 400MHz clock frequency with Innovus by Candense. The bitwidth of weight, input images data and bias data are set to 16 bits fixed point. Proposed SF-MMCN is able to execute on VGG-16 and ResNet-18 in this paper. Figure 13 is the implemented architecture of proposed SF-MMCN. According to Figure 13 there are 8 SF-MMCNs in implemented architecture.
4.1 Selection of number of SF-MMCN
Figure 15: The number of SF-MMCN and corresponded Efficiency Factor $\nu$.
As mentioned above, there are 8 proposed SF-MMCN in implemented structure. The reason why adopting 8 SF-MMCN is owing to the performance indicated in Figure 15. According to Figure 15, since efficiency factor is the ratio between power and the actual executed PE in implementation, 8 SF-MMCNs performances great $\nu$ values. Although 16 SF-MMCNs provides the best $\nu$ value, the power consumption and the number of PEs become as large as the target of this paper. In other words, 16 SF-MMCNs shows the best result of $\nu$, but performs worse efficiency (GOPs/W). On the other hands, 2 and 4 SF-MMCNs provides the unwilling value of $\nu$ because small MAC core unbalances the distribution of each hierarchies.
4.2 Results and Analysis
Figure 16: The PE utilization of Proposed SF-MMCN on (a) VGG-16. (b) ResNet-18.
Table 1: Comparison with Other Accelerators
Performance Ahmadi et al. (2021) Huang et al. (2021) Shao et al. (2021) Mo et al. (2021) Moon et al. (2023) citemy This work
@@ 4.2 Results and Analysis
CNN models VGG-16 ResNet-50 VGG-16 AlexNet VGG-16 AlexNet VGGNet GoogleNet ResNet Eff.N-L0 ViT-T M.Mxr-B VGG-16 VGG-16 ResNet-18
@@
Throughput (GOPs) 77.4/75.4 84.0 403 NA 1870-18900 2572.184 163.9
Energy Effi. (GOPs/W) 0.31k/0.3k NA 2.1k 12.1k 907k-551k 718k 9.1k
Area Effi. ($Gops/{mm^{2}}$) NA 30.55 NA 745.1 720-2600 NA 86.2
@@
4.2.1 PE utilization
The target of this paper is to reduce the scale of the PE array and increase the utilization of PEs in hardware implementation. Figure 16 (a) and Figure 16 (b) are the performance of PE utilization on VGG-16 and ResNet-18 respectively. In the first layer of VGG-16, the PE utilization is lower than any other layer owing to the input image channel is 3. Therefore, only 6 proposed SF-MMCN are set to execute the convolution. In the rest layer of the VGG-16, the PE utilization is about 89% because the series structure in VGG-16. The PE_9 in each propsoed SF-MMCNs only process data reuse function instead of MAC operations.
On ResNet-18, Figure 16 (b), the reason of PE utilization of the first layer is lower than the rest of layers is the same as VGG-16. The series structure in ResNet-18 is also about 89%. The highlight of ResNet-18 is residual structures. Due to the MAC operations and data transmission of residual block, the PE utilization is up to 100%. The definition of PE utilization in this paper is that all sub-circuits in PE are fully operated MAC computations in single cycle of convolution. Therefore, although PE utilization in the first of VGG-16 and ResNet-18 are the lowest in whole CNN models, the density of MAC operation are much higher than CARLA Ahmadi et al. (2021), which only executes 3 PEs per cycle.
4.2.2 Operation Efficiency
The PEs in every proposed SF-MMCNs are executing MAC operations in parallel. Therefore, all PEs deliver convolution outputs at the same time after 9 cycles, which means a SF-MMCN can finish 8 convolution outputs in 9 cycles. Table 2 is the operation efficiency comparison between CARLA Ahmadi et al. (2021) and proposed SF-MMCN. CARLA operates convolutions by each row of filter. Therefore, if the pixel of input feature and the filter size are 28 and $3\times 3$ respectively, CARLA has to spend around 3 times of pixel cycles to finish a convolution computation. Meanwhile, the MAC operations of proposed SF-MMCN is about ×$\times 2.67$ than CARLA due to the density of MAC operation in SF-MMCN is almost 100% in single convolution computation.
Table 2: Operation Efficiency Comparison
@@ 4.2.2 Operation Efficiency
Ahmadi et al. (2021) SF-MMCN Ahmadi et al. (2021) SF-MMCN (Normalization)
@@
4.3 Comparison with Other Accelerators
@@
The comparison of proposed SF-MMCN and other remarkable accelerators are indicated in Table 1. Although proposed SF-MMCN only implemented VGG-16 and ResNet-18, it also support other CNN models due to SF structure. This paper selects these models as examples of series and parallel models. The power and area performances are from synthesis results. First of all, as one of the targets of this study, proposed SF-MMCN only required 72 PEs. The reason why Moon et al. (2023) only adopted 8 PEs is that the specific structure of PEs in Moon et al. (2023). The PE in Moon et al. (2023) not only consists of multipliers and accumulators, but also data management circuits. What’s more, though Hsu et al. (2023) uses 32 PE in implementation, the latency of MMCN Hsu et al. (2023) on CNN models with parallel structure is not as well proposed SF-MMCN, Figure 17. Due to the quantized-MAC operations in QNAP structure, the PE structure is different from traditional PE structures. Therefore, although the PE utilization in QNAP Mo et al. (2021) is almost 100%, the density of hardware executing reports the lower hardware utilization.
@@
The throughput of each accelerators in Table 1 are the peak throughput. Though MMCN Hsu et al. (2023) is the highest among all accelerators because the different definition of throughput. The throughput OPs in this paper almost equals to FLOPs. Therefore, proposed SF-MMCN performs throughput around 163.9GOPs. Due to operating MAC computations by row per cycles in CARLA, the throughput are affected by the dataflow in Ahmadi et al. (2021) , Consequently, compared to Ahmadi et al. (2021) and Huang et al. (2021), proposed SF-MMCN achieve ×$\times 2.11$ and ×$\times 1.95$ time of which respectively.
Figure 17: Latency comparison between MMCN Hsu et al. (2023) and proposed-SF-MMCN.
@@
The other factor of performance in this paper is $\nu$. As the index of operating density of hardware, proposed SF-MMCN provide the smallest value of $\nu$ among other accelerators. Owing the specific PE structures in Huang et al. (2021), Mo et al. (2021) and Moon et al. (2023), it’s hard to receive the clear value of PE utilization ($U\_PE$ in equation (2)). Therefore, the $\nu$ in Table 1 are "-". According to Table 1, CARLA Ahmadi et al. (2021) reports the highest $\nu$ value is because only 3 PEs operates in a cycles. The $P_{act}$ only 3 when the size of filter is $3\times 3$. Consequently, the value of $\nu$ on proposed SF-MMCN is the smallest represents that proposed SF-MMCN is able to operate Neural Network computation without any redundant hardware, which reduce the power significantly.
5 Conclusion
[…]

### v2

@@ III-F Data Reuse
Leveraging the SF structure, the proposed SF-MMCN optimizes data reuse due to the prevalence of repeated input images, Figure 17 (a). Between each convolution cycle, there exist 8 repeated input data, corresponding to 8 registers from PE_1 to PE_8. To accommodate the 16-bit requirements for residual blocks in hardware, SF-MMCN expands the bit-width of these 8 registers to 32 bits, facilitating simultaneous storage of residual and reused data, Figure 17 (b). This strategic approach reduces power consumption by avoiding the reloading of repeated data in every computation cycle. As a result, SF-MMCN significantly conserves energy, demonstrating the efficacy of its SF structure.
@@
Figure 17: (a) Illustration of repeated data in the proposed SF-MMCN. (b) Data-reused structure.
III-G Implementation Architecture
@@
The implementation of the proposed SF-MMCN is indicated in Figure 18. The input image features and weight data are stored in off-chip memory before executing. After input and weight data enter the computation core through I/O interface, dataflow would be managed by TOP CTRL through weight and input buffer.
Figure 19: Dataflow comparison between traditional structures and the proposed SF-MMCN on (a) waveform. (b) CNN structure illustration.
There are four SF-MMCNs in the implementation of the proposed SF-MMCN. Owing to data reuse in implementation and large MAC in a dense layer, each SF-MMCN can exchange data by registers of each PE in SF-MMCN. If the output channel exceeds hardware limitation, the dataflow of computation will iterate to complete convolutions. What’s more, with a pooling unit and activation function unit in the proposed SF-MMCN, it also supports different functions of CNN as MMCN in [24]. All control signals are switched by TOP CTRL through the testbench.
III-H Operation Efficiency of Residual Blocks
In traditional CNN accelerators, when facing a parallel CNN structure, most of the strategy is modifying the dataflow from parallel to series [19], [24], as shown in Figure 19 (a). In this study, the proposed SF-MMCN combines a normal convolution layer and a convolution layer hidden in the residual block. The Residual_0 in figure 19 (a) doesn’t have to wait for series structure (Conv_0 and Conv_1 in Figure 19 (a)) finishing computing. In other words, according to Figure 19 (b), traditional CNN accelerators spend additional cycles to complete convolution layers with parallel structures. Therefore, the proposed SF-MMCN performs the higher efficiency while conducting parallel structures.
III-I Hardware Utilization and Power Efficiency
This paper aims to increases the utilization of PEs in PE array in a CNN accelerator. Due to the pipeline technique in a PE in the proposed SF-MMCN, a single PE can complete a convolution computation by itself. The utilization of PE in hardware is indicated in:
@@ III-I Hardware Utilization and Power Efficiency
In equation (1), Where $C_{t}$ is the percentage of computing cycles, which means actual operating cycles $T$ divided by total enable cycles $t$. $PE_{act}$ and $PE_{total}$ represent actual executing PEs and total PEs in hardware respectively. $U_{PE}$ is the utilization of PEs in hardware. Although most CNN accelerators perform high PE utilization, the occurrence of redundant PE still happens when encountering an unsuitable CNN model. The utilization of PE is relatively connected with power consumption. As the biggest structure in a CNN accelerator, the PE array consumes the most power in operation. The equation (3) indicates the total power $P_{total}$ of a CNN accelerator.
@@
N is equal to $PE_{act}$ in equation (2), which means the number of actual executing PEs in hardware. $P_{1}$ represents the power consumption of a single PE, $P_{R}$, and $P_{C}$ are the power of redundant circuit and control unit respectively. However, given the diversity of conditions in different CNN models, such as the parallel block mentioned before. What’s more, with the development of CC accelerators, many more studies perform high efficiency in implementation. $P_{total}$ only expresses the estimation of total power. Therefore, another concept of efficiency factor ($\nu$) is explained in equation (4).
@@
The efficiency factor equals the utilization of PE divided by total power consumption. According to equation (4), it reveals that if the value of $\nu$ increases, it means the corresponding structure spends much more power on redundant hardware. This is because the value of PE utilization never equals zero except the circuit is idle. Consequently, the smaller $\nu$ shows that the power consumption mostly results from the PE array. Likewise, higher PE utilization represents the well-allocated hardware, which indicates more PEs are conducting MAC operations. In other words, to reduce the power consumption of a CNN accelerator, it’s necessary to reduce the scale of the PE array, which meets the target of this paper.
Moreover, function (4) can also be computed under single convolution layers. This paper emphasizes the operation conditions of series and parallel CNN models, function (4) also provides $\nu$ under different environments by adjusting the value of $N$ and $P_{R}$ in equation (4).
IV Experimental results
The proposed SF-MMCN is implemented in Verilog HDL and logic synthesis with Design Compiler by Synopsys and implemented in TSMC 40nm technology under 400MHz clock frequency with Innovus by Candense. The bit-width of weight, input images data, and bias data are set to 16 bits fixed point. The proposed SF-MMCN can execute on VGG-16 and ResNet-18 in this paper. Figure 18 is the implemented architecture of the proposed SF-MMCN. According to Figure 18 there are 8 SF-MMCNs in implemented architecture.
IV-A Selection of number of SF-MMCN
Figure 20: The number of SF-MMCN and corresponding Efficiency Factor $\nu$.
As mentioned above, there are 8 proposed SF-MMCN in implemented structure. The reason why adopting 8 SF-MMCN is owing to the performance indicated in Figure 20. According to Figure 20, since the efficiency factor is the ratio between power and the actual executed PE in implementation, 8 SF-MMCNs perform great $\nu$ values. Although 16 SF-MMCNs provide the best $\nu$ value, the power consumption and the number of PEs become as large as the target of this paper. In other words, 16 SF-MMCNs shows the best result of $\nu$, but perform worse efficiency (GOPs/W). On the other hand, 2 and 4 SF-MMCNs provide the unwilling value of $\nu$ because a small MAC core unbalances the distribution of each hierarchy.
IV-B Results and Analysis
Figure 21: The PE utilization of the proposed SF-MMCN on (a) VGG-16. (b) ResNet-18.
TABLE I: Comparison with Other Accelerators
Performance TCASI’21[15] TCASI’21[28] TCASI’22[29] ISSCC’21[19] ISSCC’23[30] MMCN[24] This work
@@ IV-B Results and Analysis
CNN models VGG-16 ResNet-50 VGG-16 AlexNet VGG-16 AlexNet/VGGNet GoogleNet/ResNet Eff.N-L0 ViT-T/M.Mxr-B VGG-16 VGG-16 ResNet-18
@@
Throughput (GOPs) 77.4/75.4 84.0 403 NA 1870-18900 2572.184 437.9
Energy Effi. (GOPs/W) 0.31k/0.3k NA 2.1k 12.1k 907k-551k 718k 24.3k
Area Effi. ($Gops/{mm^{2}}$) 12.48 30.55 NA 745.1 720-2600 NA 230.47
@@
IV-B1 PE utilization
The target of this paper is to reduce the scale of the PE array and increase the utilization of PEs in hardware implementation. Figure 21 (a) and Figure 21 (b) are the performance of PE utilization on VGG-16 and ResNet-18 respectively. In the first layer of VGG-16, the PE utilization is lower than any other layer owing to the input image channel being 3. Therefore, only 6 of the proposed SF-MMCN are set to execute the convolution. In the rest layer of the VGG-16, the PE utilization is about 89% because of the series structure in VGG-16. The PE_9 in each of the proposed SF-MMCNs only processes the data reuse function instead of MAC operations.
On ResNet-18, Figure 21 (b), the reason for PE utilization of the first layer is lower than the rest of the layers is the same as VGG-16. The series structure in ResNet-18 is also about 89%. The highlight of ResNet-18 is residual structures. Due to the MAC operations and data transmission of the residual block, the PE utilization is up to 100%. The definition of PE utilization in this paper is that all sub-circuits in PE are fully operated MAC computations in a single cycle of convolution. Therefore, although PE utilization in the first of VGG-16 and ResNet-18 are the lowest in whole CNN models, the density of MAC operation are much higher than CARLA [15], which only executes 3 PEs per cycle.
TABLE II: Operation Efficiency Comparison
@@ IV-B1 PE utilization
[15] SF-MMCN [15] SF-MMCN (Normalization)
@@
IV-B2 Operation Efficiency
The PEs in every proposed SF-MMCN are executing MAC operations in parallel. Therefore, all PEs deliver convolution outputs at the same time after 9 cycles, which means an SF-MMCN can finish 8 convolution outputs in 9 cycles. Table II is the operation efficiency comparison between CARLA [15] and the proposed SF-MMCN. CARLA operates convolutions by each row of filters. Therefore, if the pixel of the input feature and the filter size are 28 and $3\times 3$ respectively, CARLA has to spend around 3 times of pixel cycles to finish a convolution computation. Meanwhile, the MAC operations of the proposed SF-MMCN is about ×$\times 2.67$ than CARLA due to the density of MAC operation in SF-MMCN is almost 100% in single convolution computation.
Figure 22: The MAC operation performance of the proposed SF-MMCN.
To analyze the operation efficiency of the proposed SF-MMCN, Figure 22 and Figure 23 indicate the performance of the proposed structure under same and different pixels of input feature map and weight data. When computing under the same size of weight data, it’s apparent that the number of cycles to deliver the first MAC output of the proposed SF-MMCN maintains only 9 despite the increasing of input data Figure 22. On the other hand, CARLA [15] requires additional cycles due to a modified low-power structure. The smallest cycle for CARLA has always been 3 times N according to Figure 22. Once the size of input data exceeds 32, CARLA [15] consumes many more cycles finishing convolution.
However, the size of weights is not always the same owing to the more complex CNN models coming out. Therefore, the analysis of the corresponding circumstance is revealed in Figure 23. Where Wh and Ww represent the height and width of weight data respectively. The result shows that though the proposed structure requires to complete whole weight pixel in one convolution operation, the self-computing characteristic of single PE makes SF-MMCN deliver 9 convolution outputs in these cycles. However, CARLA [15] only provides one convolution output in the same cycle on account of computing input image per row. Consequently, combined with Table II and Figure 23, the result shows that the proposed SF-MMCN can keep in high operation efficiency under different environments.
Figure 23: The efficiency performance of the proposed SF-MMCN.
Figure 24: Latency comparison between MMCN [24] and the proposed-SF-MMCN.
The highlight of the implementation of the proposed SF-MMCN on diffusion model is the throughput performance. Since the complexity of a single block in U-net is much more than a single convolution layer in VGG or ResNet. Therefore, the best case of throughput is Block 2 and Block 3 according to Figure 14 and Table 25. These are two of the most massive layers in a block in U-net. Since Block 1 and Block 4 are only for single channels or simple computation, which result in light percentages of operation cycles in one block in U-net. Therefore, if considering 4 blocks together to observe the throughput of a single block, it will achieve up to 437.976 GOPs. This result indicates that SF-MMCN successfully supports the diffusion model with high efficiency.
Figure 25: The throughput performance of the proposed SF-MMCN on U-net.
IV-C Comparison with Other Accelerators
@@
The comparison of the proposed SF-MMCN and other remarkable accelerators is indicated in Table I. Although the proposed SF-MMCN only implemented VGG-16 and ResNet-18, it also supports other CNN models due to SF structure. This paper selects these models as examples of series and parallel models. The power and area performances are from synthesis results. First of all, as one of the targets of this study, the proposed SF-MMCN only required 72 PEs. The reason why [30] only adopted 8 PEs is that the specific structure of PEs in [30]. The PE in [30] not only consists of multipliers and accumulators but also data management circuits. What’s more, though [24] uses 32 PE in implementation, the latency of MMCN [24] on CNN models with parallel structure is not as well as the proposed SF-MMCN, Figure 24. Due to the quantified MAC operations in the QNAP structure, the PE structure is different from traditional PE structures. Therefore, although the PE utilization in QNAP [19] is almost 100%, the density of hardware executing reports lower hardware utilization.
@@
The throughput of each accelerator in Table I is the peak throughput. Though MMCN [24] is the highest among all accelerators because of the different definitions of throughput. The throughput OPs in this paper are almost equal to FLOPs. Therefore, the proposed SF-MMCN performs a throughput of around 437.9 GOPs. Due to a large number of parameters and multiple layers in U-net, the proposed SF-MMCN achieves throughput incredibly compared to the old version MMCN [24].
With high throughput performances of the proposed design, the other representative FoMs are operation efficiency $GOPs/W$ and area efficiency $GOPs/{mm^{2}}$. The former indicates the performance of the throughput of every power consumption unit of the accelerator, the latter is introduced in terms of the ratio of the throughput and area. According to Table I, all the information of the other state-of-the-art CNN accelerators is the result shown in corresponding references. Consequently, the operation efficiency of the proposed design is up to 81 times compared to CARLA [15]. In area efficiency, the proposed structure advances nearly by 18 times (18.42).
@@
The other factor of performance in this paper is $\nu$. As the index of operating density of hardware, the proposed SF-MMCN provides the smallest value of $\nu$ among other accelerators. Owing the specific PE structures in [28], [19] and [30], it’s hard to receive the clear value of PE utilization ($U\_PE$ in equation (2)). Therefore, the $\nu$ in Table I are ”-”. According to Table I, CARLA [15] reports the highest $\nu$ value is because only 3 PEs operates in cycles. The $P_{act}$ only 3 when the size of the filter is $3\times 3$. Consequently, the value of $\nu$ on the proposed SF-MMCN is the smallest representing that the proposed SF-MMCN can operate Neural Network computation without any redundant hardware, which reduces the power significantly.
Figure 26: The layout of the proposed SF-MMCN chip.
IV-D Final Implementation
TABLE III: Performance of the proposed SF-MMCN Chip
Performance Specifications
Technology TSMC 40 nm
Frequency 200 MHz
Voltage 0.9 V
Bit-width 16 bits
Chip size 0.63mm $\times$ 0.63mm
Chip area (Core) 0.39 $mm^{2}$
Total Power 116.7 mW
Efficiency 3.75 $GOPs/mW$
Area Effi. 3752.36 $GOPs/{mm^{2}}$
The layout depicted in Figure 26 represents the proposed SF-MMCN chip design. Upon completing the placement and routing processes, the final performance metrics of the chip are summarized in Table III. Notably, the proposed SF-MMCN chip exhibits remarkable efficiency, reaching up to 3.75 ${GOPs}/{mW}$ while occupying a mere 0.17 ${mm^{2}}$ of hardware area. The throughput achieved is attributed to the operation of eight SF-MMCN cores. Scaling up the implementation by incorporating additional SF-MMCN cores promises a significant boost in throughput, facilitated by self-convolution by a processing element (PE). Furthermore, the compact hardware footprint suggests the potential for scalability without incurring substantial increases in hardware and power consumption. In conclusion, the proposed SF-MMCN chip successfully meets the objectives outlined in the thesis.
V Conclusion
This paper provides the structure of SF-MMCN, a CNN accelerator with SF structure which lets SF-MMCN perform high throughput and acceleration of not only series but also parallel CNN structure. SF structure in the proposed SF-MMCN can operate convolution and conduct data reuse. SF-MMCN performs 18 mW of power consumption with only 1.9 $mm^{2}$ hardware area. The number of PE in SF-MMCN implementation is only 72, which is lower than other accelerators with traditional PE structures. The concept of efficient factor $\nu$ is also proposed in this paper, which represents the utilization density of PEs in an accelerator. Finally, SF-MMCN achieves energy efficiency 24.3 k($GOPs/W$) and 230.47($GOPS/mm^{2}$) of area efficiency.